Define the terms: elastic limit, yield point, elastic fatigue, and breaking stress. Establish a relationship among Young’s modulus (Y), bulk modulus (K), rigidity modulus (η), and Poisson’s ratio (σ). 


1
Expert's answer
2021-03-09T15:30:41-0500

Elastic limit is the upper limit of deforming force upto which, if deforming force is removed, the body regains its original form completely and beyond which if deforming force is increased the body loses its property of elasticity and get permanently deformed.

The yield point is the point on a stress-strain curve that indicates the limit of elastic behavior and the beginning of plastic behavior.

Elastic fatigue is a property of the elastic body by virtue of which it's beheviour becomes elastic under the action of repeated alternating deforming force.

Breaking stress is the maximum force that can be applied on a cross sectional area of a material in such a way that the material is unable to withstand any additional amount of stress before breaking.


"\frac{3}{Y}=\frac{1}{\eta}+\frac{1}{3K}\\2\sigma=1-\frac{Y}{3K}"